OpenStudy (anonymous):

[1.03] Evaluate the following expression using the values given. Find 2x − 3y − z if x = −3, y = 2, and z = −2

OpenStudy (anonymous):

All you have to do is replace the variables with numbers because you already know them!

OpenStudy (anonymous):

_10?

OpenStudy (anonymous):

Yes! The answer is negative ten!

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

2(-3) -3(2) -(-2) = -6 -6 + 2 = -10

OpenStudy (anonymous):

If you're satisfied with the answers to your question, when you're done you should make sure to close it :)